OpenStudy (anonymous):

what is the graph of the function 1/2 x^2

OpenStudy (gorv):

\[\frac{ x^2 }{ 2 }\]

OpenStudy (gorv):

it is ??

OpenStudy (gorv):

@princessbae98

OpenStudy (anonymous):

ok I got the graphs I need but im still confused

OpenStudy (gorv):

put x=2

OpenStudy (gorv):

\[y=\frac{ x^2 }{ 2 }\]

OpenStudy (anonymous):

ok but then what would the graph be if x=2

OpenStudy (gorv):

at x=0 y=0 so it will start from origin

OpenStudy (gorv):

just plug in the value x=2

OpenStudy (gorv):

\[y=\frac{ 2^2 }{ 2 }\]

OpenStudy (gorv):

2^2=??

OpenStudy (anonymous):

4

OpenStudy (gorv):

4/2=???

OpenStudy (anonymous):

2

OpenStudy (gorv):

y=2

OpenStudy (anonymous):

ok thank you

OpenStudy (gorv):

check in graph when x=2 y=2

OpenStudy (anonymous):

I found the answer thanks can you help me with another question

OpenStudy (gorv):

yeah post new Q